OpenStudy (anonymous):

2(10^x)=14? help solve equation

OpenStudy (tkhunny):

Have you considered dividing by 2 and introducing a logarithm?

OpenStudy (anonymous):

No I haven't missed a lot of class recently but thanks for pointing me in the right direction!

OpenStudy (anonymous):

2*10=20. plug that in: 20^x=14 divide by 2 10^x=7 and if i am not mistaken the log would be: log(10)7=x......@tkhunny correct me if im wrong

Parth (parthkohli):

Uh, the answer is right, but the method is pretty wrong.

Parth (parthkohli):

I mean, you should divide both sides by 2 and then introduce that logarithm.

OpenStudy (tkhunny):

No, "2*10" is not appropriate for this problem, Oddly, though, 10^x = 7 is correct.

OpenStudy (anonymous):

correct it up haha. its been a while since ive done that. based on what youre saying, i kinda took the round about way of doing it. well @tkhunny basically the 2*10 is doing the same thing, because the whole thing is eventually divided by 2 in the end. like i said i just kinda took the round about way of doing it

Parth (parthkohli):

...Eh?

OpenStudy (tkhunny):

No. That is just no good. It is only by luck that you ended up in the right place.

Parth (parthkohli):

So you're saying that \(2 \times 5^3\) is \(\dfrac{10^3}{2}\)

OpenStudy (anonymous):

@ParthKohli huh? where is that 5^3 coming from?

Parth (parthkohli):

I was just giving an example why your method is fallacious.

Parth (parthkohli):

:-)

OpenStudy (anonymous):

haha ok, welll i will trust you then, i am bowing out my friends